A certain alloy is a mixture of equal number A atoms and B atoms. B atoms are twice as heavy as A atoms. What weight of A atoms is needed to make 600 pounds of the alloy?
Answer #1

Given that number of atoms are equal, So

Na = Nb

Also, B atoms are twice as heavy as A atoms

Wb = 2*Wa

Total weight of alloy = 600 Pound

Wa + Wb = 600 pounds

Since Wb = 2*Wa, So

Wa + 2*Wa = 600

3*Wa = 600

Wa = Weight of A atoms = 200 pounds

Please Upvote.

Are you sure this is complete question? Please check your question again if above answer does not work and let me know if there is anything missing in the above question.